For a segment of a radio show , a disk jockey can play nine records. If there are 12 records to select from, in how many ways can the program for this segment be arranged?
Answer by fractalier(6550)

12 chosen 9 at a time is calculated by
12! / (12-9)!9! = 12*11*10 / 3*2*1 = 220 ways